OpenStudy (anonymous):

(x^2)/(x^(4)-81) find the horizontal, vertical, and oblique asymptotes

OpenStudy (asnaseer):

ok, first lets find the vertical asymptotes - do you know how to find these?

OpenStudy (anonymous):

no i dont

OpenStudy (asnaseer):

sorry - ignore that - a vertical asymptote is found by finding a value for x that makes y infinite

OpenStudy (asnaseer):

that generally means finding a value for x that will make the denominator of a rational polynomial (i.e. a polynomial in terms of a fraction) is to zero

OpenStudy (asnaseer):

can you see what value of x would make the denominator zero?

OpenStudy (asnaseer):

i.e. we want to find values for x such that:\[x^4-81=0\]

OpenStudy (anonymous):

9?

OpenStudy (asnaseer):

\(9^2=81\) so this doesn't work

OpenStudy (asnaseer):

you have a 4'th power here

OpenStudy (anonymous):

oh so 3?

OpenStudy (asnaseer):

almost: \(3^4=81\) yes, but there is also another value - can see what that would be?

OpenStudy (asnaseer):

hint: \(2^2=4\) and \((-2)^2=4\)

OpenStudy (anonymous):

-3

OpenStudy (asnaseer):

perfect! so now we know the two vertical asymptotes, they are x=3 and x=-3

OpenStudy (asnaseer):

the horizontal and/or oblique asymptotes are found by looking at what the given polynomial will "tend" to as x tends to plu/minus infinity

OpenStudy (asnaseer):

so, in this case we have:\[y=\frac{x^2}{x^4-81}\]now think about what the denominator would tend to as x tends to infinity - can you see that the "- 81" term will become more and more insignificant as x tends to infinity?

OpenStudy (anonymous):

yeah

OpenStudy (asnaseer):

good, so we can express this as:\[\lim_{x\rightarrow\infty}\frac{x^2}{x^4-81}=\frac{x^2}{x^4}=\frac{1}{x^2}\]does that make sense?

OpenStudy (anonymous):

ya kinda

OpenStudy (asnaseer):

ok - let me know if you need more clarification. now also notice that since we have even powers in the numerator AND the denominator, the the limit will be the same for x tending to minus infinity - agreed?

OpenStudy (asnaseer):

*then the

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

good: we can express this as follows:\[\lim_{x\rightarrow\pm\infty}\frac{x^2}{x^4-81}=\frac{x^2}{x^4}=\frac{1}{x^2}\]

OpenStudy (asnaseer):

now, what do you think the value of \(\displaystyle\frac{1}{x^2}\) will be as x tends to plus/minus infinity?

OpenStudy (asnaseer):

imagine x getting bigger and bigger - what will happen to the value of \(\displaystyle\frac{1}{x^2}\) in this case?

OpenStudy (anonymous):

i have no idea.

OpenStudy (anonymous):

I'm sry i am so horrible at math

OpenStudy (asnaseer):

np - let me try and explain it....

OpenStudy (asnaseer):

lets take different values of x: x=10 ---> 1/x^2 = 1/100 = 0.01 x=100 ---> 1/x^2 = 1/10000 = 0.0001 ... can you see the value of 1/x^2 gets smaller and smaller as x gets bigger and bigger?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good, so eventually, when x is plus/minus infinity, 1/x^2 will become zero

OpenStudy (asnaseer):

so we have shown that as x tends to plus/minus infinity, the value of \(\displaystyle\frac{x^2}{x^4-81}\) tends to zero. this means we have a horizontal asymptote of y=0.

OpenStudy (asnaseer):

you can see these asymptotes here: http://www.wolframalpha.com/input/?i=asymptotes+%28x^2%29%2F%28x^%284%29-81%29

OpenStudy (asnaseer):

let me know if you are still unsure of anything here.

OpenStudy (anonymous):

cool thank you so much!

OpenStudy (asnaseer):

yw :)